Let X and Y be two topological spaces with Y Hausdorff, and let f,g : X → Y be two continuous maps. Show that: (a) If f is one-to-one, then X is Hausdorff; (b) If f(x) = g(x) ∀x ∈ A, where ¯A = X, then f = g; (c) The graph Γ = { (x,f(x)) | x ∈ X} is closed in X ×Y .


1
Expert's answer
2021-03-23T06:21:22-0400

"a)\ \text{We take two elements } x_1\text{ and } x_2\in X, x_1\neq x_2.\\
\text{Then we find }f(x_1) \text{ and } f(x_2), f(x_1)\neq f(x_2).\\
\text{As } Y \text{ is Hausdorff there exist two neighbourhoods }U\text{ and }\\
V\text{ in Y}: f(x_1)\subset U,\ f(x_2)\subset V,\ U\cap V=\empty.\\
\empty=f^{-1}(U\cap V)=f^{-1}(U)\cap f^{-1}(V).\\
\text{We found two neighbourhoods }f^{-1}(U)\text{ and }f^{-1}(V)\\
\text{of } x_1 \text{ and } x_2\text{ respectively such that }f^{-1}(U)\cap f^{-1}(V)=\empty.\\
X\text{ is Hausdorff}.\\
b)\text{ We should show that } f(x)=g(x)\text{ on the set of all the boundary points of }X.\\
\text{Let }x_0\text{ be a boundary point of }X.\\
\text{We reach }x_0\text{ from the side of }X:\\
\lim_{x\rightarrow x_0}f(x)=f(x_0)\text{ (f is continuous)}.\\
\lim_{x\rightarrow x_0}g(x)=g(x_0)\text{ (g is continuous)}.\\
f(x)=g(x)\text{ for all the interior points. So }f(x_0)=g(x_0)."

c) We take a point "(x,y)\notin \Gamma." Because "y\neq f(x)" and "Y" is Hausdorff, there exist disjoint open sets "V,\ V^{\prime}" with "f(x)\in V" and "y\in V^{\prime}". "f" is continuous, so "U=f^{-1}(V)" is open with "x\in U". We see that "U\times V^{\prime}" is an open neighborhood of "(x,y)" that does not intersect the graph. Indeed, if we suppose (for contradiction) that "(z,f(z))\in U\times V^{\prime}", then it would follow that "f(z)\in V^{\prime}". However, "z\in U=f^{-1}(V)" which means that "f(z)\in V". Thus, "f(z)∈V\cap V^{\prime}" contradicting the fact that "V,\ V^{\prime}" are disjoint.